OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

ganeshie8 (ganeshie8):

is that \(7\sin^2 x\) ?

OpenStudy (anonymous):

is the question correct?

OpenStudy (anonymous):

Yes it is (7sin^2x)

ganeshie8 (ganeshie8):

add 5 both sides

ganeshie8 (ganeshie8):

then divide 7

OpenStudy (anonymous):

Then, substitute sin x = t, solve the quadratic. Check whether the solutions lie in the range of sine fn.

OpenStudy (anonymous):

To get sin^2x-2sinx+1? Ok then to solve the quadratic I would make it (sin +2)(sin-1)

ganeshie8 (ganeshie8):

u shoudl get (sinx -1)^2

ganeshie8 (ganeshie8):

check once

OpenStudy (anonymous):

Oh ok I see. . . Where do i go from there?

ganeshie8 (ganeshie8):

(sinx -1)^2 = 0

ganeshie8 (ganeshie8):

solve for sinx, then solve x

OpenStudy (anonymous):

What would be my first step?

ganeshie8 (ganeshie8):

take square root (sinx -1)^2 = 0 sinx-1 = 0

ganeshie8 (ganeshie8):

sinx = 1

OpenStudy (anonymous):

Wow that's great. I need to find all of the solutions now. How would I go about that?

ganeshie8 (ganeshie8):

when does sinx equals 1 ?

ganeshie8 (ganeshie8):

when x = pi/2 right ?

ganeshie8 (ganeshie8):

and since sin period is 2pi, for every 2pi, it gets the same value

OpenStudy (anonymous):

Yes. Oh ok I like where this is going. I can find the other solutions knowing pi/2.

OpenStudy (anonymous):

so 2pi/3, 4pi/3, 5pi/3

ganeshie8 (ganeshie8):

good :) it wud be simply \(\pi/2 + 2n \pi\)

OpenStudy (anonymous):

wait no no no

ganeshie8 (ganeshie8):

n is any integer

ganeshie8 (ganeshie8):

ha

OpenStudy (anonymous):

OK thank you so much

ganeshie8 (ganeshie8):

np :)